How many C atoms are contained in a 2.5 mol of C2H3O2F? Show factor unit method conversion please
Answer #1

One mole of the compound C2H3O2F has 2 moles of carbon atom per formula unit of the compound

Therefore, 2.5 moles of the compound will have = 2×2.5 = 5 moles

Therefore, number of carbon atom = moles ×NA = 5×6.022×1023 = 3.011×1024 atoms